Suppose a Toyota C-HR car, sold in Europe, was designed to run on a mix of ethanol and gasoline. Suppose that ethanol can be manufactured in an ideal situation for a cost of approximately $0.85 per gallon. Therefore, when blended with gasoline costing $6.00 per gallon, a mixture of 15% ethanol and 85% gasoline (called E15) costs $5.23 per gallon. How much money would be saved if you drive 12,000 miles in one year using the E15 fuel instead of pure gasoline? Assume fuel consumption at 30 miles per gallon in both cases, and assume that your engine's performance is not adversely affected by using the E15 blend instead of pure gasoline. Provide your answer to the nearest dollar.
Answer #1

Fuel consumption=total miles traveled/Miles traveled with 1 gallon=12000/30= 400 gallons

Fuel expenditure with E15=Price per gallon*Fuel consumption=5.23*400=$2092

Fuel expenditure with Pure gasoline=Price per gallon*Fuel consumption=6*400=$2400

Saving with E15 fuel=2400-2092 =$308